An automobile makes a turn whose radius 18 m. The road is banked at an angle of

10o when the friction between the tires and the road is zero. At what speed should the driver take the curve to avoid sliding off the road?


1
Expert's answer
2021-01-16T17:22:10-0500

"N\sin10°=m\frac{v^2}{r}"


"N\cos10°-mg=0"


So, we have


"\tan10°=\frac{v^2}{gr}"


"v=\sqrt{9.8\cdot\tan10°\cdot 18}=5.58(m/s)" . Answer